OpenStudy (anonymous):

Write in logarithmic form. (a) e^0=1 (b) 5^x=6 (c) e^3t=8

zepdrix (zepdrix):

\[\Large a^b=c\qquad\to\qquad \log_a c=b\]

zepdrix (zepdrix):

So for the first one:\[\Large 5^x=6\qquad\to\qquad \log_5 6=x\]

OpenStudy (anonymous):

How about e^0=1? I don't understand 'e'.

zepdrix (zepdrix):

Oh I just did the second one :) lol. woops

zepdrix (zepdrix):

\[\Large e\approx2.71828\]It's just a fancy number like pi ^^ We would still follow the same format as before:\[\Large e^0=1\qquad\to\qquad \log_e 1=0\]Since the log of base e is used quite often, we have a special notation for it. We use the natural log (ln instead of log) to show that it's a log of base e:\[\Large \to\qquad \ln 1=0\]

OpenStudy (anonymous):

Ohhh. I remember now. :D

zepdrix (zepdrix):

Personally I find going from logarithmic form to exponential easier to understand. This is my thought process, maybe it'll help :)